An infinite line of charge produces a field of magnitude 5.4 × 104 N/C at a distance of 2.4 m. Calculate the linear charge density.
Answer #1

from Gauss law we have

E = 2kx/r

where x=linear charge density,

r=distance from the wire=2.4 m

k =9 x 10^9 SI units.

we have been given E = 5.4x 10^4 N/C

so plugging numbers we get

E = 2kx/r

5.4x 10^ 4 = 2*9 x 10^9*x/2.4

5.4x 10^ 4 *2.4 =2*9 x*10^9x

x =5.4x 10^ 4 *2.4 /2*9 x*10^9

x = 7.2 x 10^ -6 C/ M (coulomb per meter) =linear charge density